Depakote is a medication whose purpose is to reduce the pain associated with migraine headaches. In clinical trials of Depakote, 2% of the patients in the study experienced weight gain as a side effect. Suppose a random sample of 30 Depakote users is obtained. Find the following probabilities:
a. Exactly 3 experienced weight gain as a side effect
b. 3 or fewer experienced weight gain as a side effect
c. 4 or more experienced weight gain as a side effect.
d. Between 1 and 4 patients, exclusive, experienced weight gain as a side effect.

Answer by stanbon(75887)

Depakote is a medication whose purpose is to reduce the pain associated with migraine headaches.
In clinical trials of Depakote, 2% of the patients in the study experienced weight gain as a side effect.
Suppose a random sample of 30 Depakote users is obtained.
----
Binomial Problem with n = 30 and p(side effect) = 0.02
--------------------
Find the following probabilities:
a. Exactly 3 experienced weight gain as a side effect
P(x = 3) = 30C3*0.02^3*0.98^27 = 0.0188
-------------------------------------------
b. 3 or fewer experienced weight gain as a side effect
P(0<= x <=3) = binomcdf(30,0.02,3) = 0.9971
----------------------------------------------
c. 4 or more experienced weight gain as a side effect.
P(4<= x <=30) = 1 - P(0<= x <=3) = 0.0029
-----------------------------------
d. Between 1 and 4 patients, exclusive, experienced weight gain as a side effect.
P(1<= x <=4) = binomcdf(30,0.02,4) - binomcdf(30,0.02,0) = 0.4542
==================
Cheers,
Stan H.
==================